OpenStudy (anonymous):

What does 5 and 6/8 equal?

OpenStudy (anonymous):

What's your question?

OpenStudy (anonymous):

5.75

OpenStudy (anonymous):

What does 5 and 6/8 equal?

OpenStudy (anonymous):

Does that answer your question?

OpenStudy (anonymous):

5 6/8 = 5.75 5 6/8 = 46/8

OpenStudy (anonymous):

thk u

OpenStudy (anonymous):

Medal? :)